# LatentExplainer: Explaining Latent Representations in Deep Generative Models with Multimodal Large Language Models

### v5

@@ Appendix B Uncertainty-aware Implementations
We follow Zhu et al. [2024] to denote the true label of the interpretability for the i-th latent variable $y_{i}$ as 1 if at least two of the three annotators can see a clear pattern in the generated images, otherwise we denote it as 0. The certainty score of the explanation for the i-th latent variable is $\bar{s}_{i}$. We adopt the Jaccard Index to measure similarity between the predicted label and the true label,
@@
$\begin{split}&E(\varepsilon)=\texttt{Jaccard}(f(\bar{s}_{i},\varepsilon),y_{i}),\text{ where}\\ &f(\bar{s}_{i},\varepsilon)=\mathbbm{1}(\bar{s}_{i}\geq\varepsilon).\end{split}$
@@
By solving this equation across all datasets, we find the threshold $\varepsilon=0.2617$. Since $\varepsilon$ is solved across all datasets, it reduces the dependence on any specific dataset. Our final output $\hat{r}$ is the response that has the highest mean pairwise cosine similarity with other responses if the certainty score is equal or greater than the threshold $\varepsilon$. Otherwise, $\hat{r}$ will be “no clear explanation”.
Appendix C Human Annotaions
@@
The ground-truth annotations of explanations are performed by four annotators from the United States and China. All annotators are students that had at least an undergraduate degree. Annotators were presented with the same images and the prompts as MLLMs and were asked to annotate the pattern of the images. If there is no clear pattern, just write “No clear explanation”. The annotations are then aggregated as references to calculate the automated evaluation metrics.
@@
Appendix D An example of automatically generating prompts from an inductive bias formula
@@
Step 1: Extracting mathematical symbols: Extracting all mathematical symbols from the given formula “$p\left(z_{i}\mid t=t_{1}\right)\neq p\left(z_{i}\mid t=t_{2}\right),\forall t_{1}\neq t_{2}"$. Symbols extracted are $\mathcal{S}=\left[\right.$$p\left(z_{i}\mid\cdot\right)"$, “$\left.t=t_{1},t=t_{2},\forall t_{1}\neq t_{2}",``\neq"\right]$.
@@
Appendix E More Qualitative Evaluation Results
We provide more qualitative evaluation results: (1) the disentanglement bias: DDPM in Figure 5, and $\beta$-TCVAE in Figure 6; (2) the combination bias: CAVAE in Figure 7; (3) the conditional bias: Stable Diffusion in Figure 8 and CSVAE in Figure 9.
Figure 5: Example of generated explanations for DDPM under the disentanglement bias.
Figure 6: Example of generated explanations for $\beta$-TCVAE under the disentanglement bias.
Figure 7: Example of generated explanations for CSVAE under the combination bias.
Figure 8: Example of generated explanations for Stable Diffusion under conditional bias.
Figure 9: Example of generated explanations for CSVAE under conditional bias.
Figure 10: Visualization of the generated explanations with the inductive bias prompt for the combination bias.
Figure 11: Visualization of the generated explanations with the inductive bias prompt for the conditional bias.
Figure 12: Visualization of the generated explanations without the inductive bias prompt for the conditional bias.
Appendix F Visualization of the Generated Explanations
In this section, we provide visualization examples of the explanations generated by our proposed LatentExplainer framework under different inductive biases: combination bias with the inductive bias prompt in Figure 10, conditional bias with the inductive bias prompt in Figure 11 and without the inductive bias prompt Figure 12. Each figure demonstrates the patterns of change in generated image sequences along specific latent directions, accompanied by the corresponding explanations.
Appendix G Case Study
@@
Table 4: Examples of automatically generated inductive bias prompts

### v6

@@ 5.5. Uncertainty-aware Implementations
We follow (Zhu et al., 2024) to denote the true label of the interpretability for the i-th latent variable $y_{i}$ as 1 if at least two of the three annotators can see a clear pattern in the generated images, otherwise we denote it as 0. The certainty score of the explanation for the i-th latent variable is $\bar{s}_{i}$. We adopt the Jaccard Index to measure similarity between the predicted label and the true label,
@@
$\begin{split}&E(\varepsilon)=\texttt{Jaccard}(f(\bar{s}_{i},\varepsilon),y_{i}),\text{ where}\\ &f(\bar{s}_{i},\varepsilon)=\mathds{1}(\bar{s}_{i}\geq\varepsilon).\end{split}$
@@
By solving this equation across all datasets, we find the threshold $\varepsilon=0.2617$. Since $\varepsilon$ is solved across all datasets, it reduces the dependence on any specific dataset. Our final output $\hat{r}$ is the response that has the highest mean pairwise cosine similarity with other responses if the certainty score is equal or greater than the threshold $\varepsilon$. Otherwise, $\hat{r}$ will be ”no clear explanation”.
5.6. Human Annotaions
@@
The ground-truth annotations of explanations are performed by three annotators from the United States and China. All annotators are students that had at least an undergraduate degree. Annotators were presented with the same images and the prompts as MLLMs and were asked to annotate the pattern of the images. If there is no clear pattern, just write ”No clear explanation”. The annotations are then aggregated as references to calculate the automated evaluation metrics.
5.7. Quantitative Analysis
For the quantitative explanation evaluation, we use BLEU (Papineni et al., 2002), ROUGE-L (Lin, 2004), SPICE (Anderson et al., 2016), BERTScore(Zhang et al., 2019), and BARTScore (Yuan et al., 2021) as the automated metrics to assess the generated explanations. BLEU, and ROUGE-L are n-gram-based metrics that measure the overlap between generated and reference texts. SPICE compares scene graphs derived from the generated and reference texts. BERTScore and BARTScore utilize pre-trained transformer-based language models to compute contextual embeddings of the generated and reference texts. These metrics together provide a comprehensive assessment of both the lexical and semantic quality of the explanations.
We evaluate the performance of Gemini 1.5 Pro, Claude 3.5 Sonnet and GPT-4o across all diffusion and vae models and datasets. Our proposed LatentExlainer method is a model-agnostic module and the inclusion of LatentExplainer consistently leads to substantial performance improvements on all MLLMs – Gemini 1.5 Pro, Claude 3.5 Sonnet and GPT-4o. Across diffusion-based models in Table 2, LatentExplainer delivers significant improvements across all datasets and metrics. Notably, when integrated with GPT-4o on CelebA-HQ, BLEU improves from 5.79 to 18.50, ROUGE-L from 23.89 to 40.85, and SPICE from 12.68 to 23.90, highlighting nearly 2x gains. Similar trends are observed with Claude 3.5 Sonnet for DDPM, where BLEU on AFHQ increases from 4.16 to 12.52, and ROUGE-L rises from 22.25 to 31.68, and SPICE from 13.34 to 20.05. Even for weaker baselines like Gemini 1.5 Pro, LatentExplainer consistently improves both lexical and semantic quality, demonstrating its effectiveness in enhancing explainability for latent variables.
For VAE-based models in Table 3, LatentExplainer achieves even more striking gains. On the $\beta$-TCVAE model with GPT-4o, performance on 3DShapes sees BLEU increase from 5.51 to 25.40, and SPICE from 7.92 to 22.78. Across dSprites, the BLEU jump from 0.00 to 12.55 and ROUGE-L from 16.19 to 37.30 further demonstrate the utility of our approach. In summary, LatentExplainer consistently boosts the interpretability and textual quality of explanations for latent variable models across settings.
5.8. Ablation Study
To understand the impact of various components in our proposed LatentExplainer, we conducted a comprehensive ablation study on GPT-4o across different models (DDPM, $\beta$-TCVAE, Stable Diffusion, and CSVAE) and inductive biases (disentanglement, combination, and conditional). Specifically, we compare the removal of inductive bias prompts (IB), the removal of uncertainty quantification (UQ), and the full model against the baseline GPT-4o. The results for each dataset are provided in Table 4 and Table 5. Removing inductive bias prompts leads to a substantial drop in all generative models. Their consistent results demonstrate that inductive bias is the most important and necessary component when explaining the latent variables of generative models. The removal of uncertainty quantification also results in a slightly decreased performance in all generative models, indicating that uncertainty quantification is also effective, though not as critical as inductive bias prompts. The full model, which incorporates both inductive bias prompts and uncertainty quantification, achieves the highest overall performance, and outperforms all baselines across all models. This confirms the necessity of both inductive bias prompts and uncertainty quantification in our LatentExplainer framework, demonstrating their significant contributions to improving explanatory performance across various generative models. It also shows that our framework can effectively design prompts for different inductive biases in generative models to improve the accuracy and reduce hallucination.
5.9. Qualitative Evaluation
To analyze the explanations for DDPM under disentanglement bias, we manipulate the latent variable along a latent direction and compare it with the one that first traverses along another latent direction and then traverses along the same latent direction. The disentangled latent variable would be invariant with respect to the variations in another latent dimension. In Figure 4, for each latent direction, we pass two image sequences and an inductive bias prompt based on the disentanglement formula to the MLLM to obtain a common latent explanation. In comparison, the explanations generated without the inductive bias prompt as shown in Figure 4 tend to show ”no clear explanation” or wrong explanations as they only align with one image sequence but do not reflect the common pattern in both image sequences in view of the inductive bias. The addition of the inductive bias prompts can assist with ruling out the variation effects of other latent variables to capture the actual meaning.
We also qualitatively evaluate the explanations for CSVAE under combination bias. We transverse a latent variable and compare it with the one that first traverses another latent variable in another group and then traverses the same latent variable, which is similar to the disentanglement bias. We then compare with the one that first transverses another latent variable in the same group and then traverses the same latent variable. As Figure 6 depicts, our model can clearly show the explanations of latent variable $z_{1},z_{2},z_{3}$ as the color of the ground, the background color, and the shape of the object. The effect of removing inductive bias prompts leads to no clear explanation like the disentanglement bias in Figure 6.
In Figure 8, we provide the ”young appearance” prompt to Stable Diffusion under conditional bias, and the explanations of all three top latent directions reflect the meaning of youth. The addition of inductive bias prompts can better identify the relation with the property of interest to capture the actual meaning of latent variables. In comparison, the one without the inductive bias prompts in Figure 8, cannot find clear explanations or simply describe the characteristics in the image sequence, lacking an abstract generalization. More qualitative evaluation results can be found in Section  6.
6. More Qualitative Evaluation Results
We provide more qualitative evaluation results for each inductive bias: (1) the disentanglement bias: DDPM in Figure 9, and $\beta$-TCVAE in Figure 10; (2) the combination bias: CSVAE in Figure 11; (3) the conditional bias: Stable Diffusion in Figure 12 and CSVAE in Figure 13.
Figure 9. Example of generated explanations for DDPM under the disentanglement bias.
7. Conclusion
8. Limitation
In this work, we use GPT-4o as one of our MLLMs to explain the latent variables. Although GPT-4o has improved vision capabilities across most tasks, we still find it incapable of identifying intricate facial features, positions, and orientations. It is necessary to enhance visual understanding to improve the explanation quality for MLLMs.
@@ 8. Limitation
Appendix A An example of automatically generating prompts from an inductive bias formula
@@
Step 1: Extracting mathematical symbols: Extracting all mathematical symbols from the given formula “$p\left(z_{i}\mid t=t_{1}\right)\neq p\left(z_{i}\mid t=t_{2}\right),$ $\forall t_{1}\neq t_{2}"$. Symbols extracted are $\mathcal{S}=\left[\right.$$p\left(z_{i}\mid\cdot\right)"$, “$t=t_{1},t=t_{2},$ $\forall t_{1}\neq t_{2}",``\neq"]$.
@@
Appendix B Case Study
@@
Figure 10. Example of generated explanations for $\beta$-TCVAE under the disentanglement bias.
Figure 11. Example of generated explanations for CSVAE under the combination bias.
Figure 12. Example of generated explanations for Stable Diffusion under conditional bias.
Figure 13. Example of generated explanations for CSVAE under conditional bias.
@@
Table 6. Examples of automatically generated inductive bias prompts